OpenStudy (anonymous):

Help for medal? Please be quick! Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 20? The graph of y = f(x) will shift left 20 units. The graph of y = f(x) will shift down 20 units. The graph of y = f(x) will shift up 20 units. The graph of y = f(x) will shift right 20 units.

OpenStudy (anonymous):

itll shift right 20 units

OpenStudy (anonymous):

Can you show me how you got that?

OpenStudy (anonymous):

sure. lemme type it out

OpenStudy (anonymous):

wait no sorry

OpenStudy (anonymous):

its up

OpenStudy (here_to_help15):

Well you know that + 20 is the y intercept right?

OpenStudy (anonymous):

because you know that + 20 is the y intercept right? That leaves with you shifting up or down. Because it is positive, which way do you think you are going to shift?

OpenStudy (here_to_help15):

That leaves with you shifting up or down. Because it is positive, which way do you think you are going to shift?

OpenStudy (anonymous):

swag @Here_to_Help15

OpenStudy (here_to_help15):

= )

OpenStudy (anonymous):

Oh, okay. So its C?

OpenStudy (anonymous):

correct

OpenStudy (here_to_help15):

yes

OpenStudy (anonymous):

Okay, thank you both!

OpenStudy (here_to_help15):

do u have another question

OpenStudy (anonymous):

Yeah, I do.

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = one thirdx + 14 thirds y = 3x + 10 y = -3x - 2 y = -one thirdx + 10 thirds

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

@Here_to_Help15 ?

OpenStudy (here_to_help15):

i will help hold a sec k

OpenStudy (here_to_help15):

i think it is just 1 sec

OpenStudy (here_to_help15):

dammit @mathstudent55 can u help plz

OpenStudy (anonymous):

He's offline I don't think he can

OpenStudy (here_to_help15):

umm @ganeshie8 help plz

OpenStudy (anonymous):

It's okay, I submitted my test and passed with a 78